# A High-Quality Dataset and Reliable Evaluation for Interleaved Image-Text Generation

### v1

@@ Abstract
Recent advancements in Large Multimodal Models (LMMs) have significantly improved multimodal understanding and generation. However, these models still struggle to generate tightly interleaved image-text outputs, primarily due to the limited scale, quality and instructional richness of current training datasets. To address this, we introduce InterSyn, a large-scale multimodal dataset constructed using our Self-Evaluation with Iterative Refinement (SEIR) method. InterSyn features multi-turn, instruction-driven dialogues with tightly interleaved image-text responses, providing rich object diversity and rigorous automated quality refinement, making it well-suited for training next-generation instruction-following LMMs. Furthermore, to address the lack of reliable evaluation tools capable of assessing interleaved multimodal outputs, we introduce SynJudge, an automatic evaluation model designed to quantitatively assess multimodal outputs along four dimensions: text content, image content, image quality, and image-text synergy. Experimental studies show that the SEIR method leads to substantially higher dataset quality compared to an otherwise identical process without refinement. Moreover, LMMs trained on InterSyn achieve uniform performance gains across all evaluation metrics, confirming InterSyn’s utility for advancing multimodal systems.
@@ 1 Introduction
Multimodal understanding and generation are critical capabilities toward artificial general intelligence. In the past two years, multimodal LLMs [1, 2, 3] have shown remarkable performance in multimodal understanding and even surpassed humans in some areas, while we have also seen many impressive advances in high-quality image generation [4, 5]. However, their restriction to generating either text or image outputs in isolation, while real-world scenarios usually require mixed-modality outputs.
@@
Recently, pioneer unified multimodal understanding and generation models (uni-MLLMS), such as Chameleon [6] and Janus-Pro [7], have shown great potential. However, they struggle to generate instruction-following interleaved image-text generation. They suffer from manifesting semantic drift, low image–text synergy, and low image quality. The main challenges are the lack of high-quality training datasets and the absence of reliable evaluation to guide model development.
Although related training datasets [8, 9, 10, 11, 12] exist, they suffer from several critical limitations:
(1) Unstable quality: Built on web-crawled sources [13, 9] or reused corpora [14, 8] with inconsistent quality and lack standardized quality control mechanisms.
(2) Limited domain coverage: Focus on narrow tasks, such as creative content generation [15], story writing [16, 17], or multimodal script writing [13], limiting their applicability to broader real-world scenarios.
(3) Low interaction complexity. The majority provide static documents, image-text pairs or single-turn prompts, failing to capture the multi-turn, deeply interleaved conversations required by real scenarios.
@@
To address the above issues, we introduce InterSyn—the first fully-automated, high-quality large-scale dataset for instruction-following, multi-turn question answering with interleaved image–text responses. InterSyn offers 1.8 M single-turn and 50,000 multi-turn dialogues spanning 8 domains, 65 categories, and 3,500 fine-grained topics. Its generation is driven by a Self-Evaluation with Iterative Refinement (SEIR) pipeline that embeds self-checking and feedback loops into every iteration, guaranteeing semantic completeness, cross-modal synergy, and contextual relevance with minimal human intervention.
Evaluation is another challenge for interleaved image-text generation. Though some related benchmarks [18, 19, 20, 21, 22] exist, there are some limitations:
(1) Limited domain and scale: small, task-specific sets cannot cover realistic multi-turn dialogue needs.
(2) Costly manual evaluation: accurate assessment still hinges on human judges, whose expense and delay hinder large-scale, rapid benchmarking.
(3) Weak alignment with human preference: current automatic metrics significantly diverge from human judgments on fine-grained multimodal reasoning.
(4) Narrow evaluation scope: emphasise surface correctness while overlooking synergy and overall answer quality.
@@
We propose SynJudge, a reliable and comprehensive judge model for interleaved image-text generation, which aligns with human judgment. SynJudge can provide interpretable and quantitative feedback, facilitating more effective model training and refinement. The interleaved image-text response is assessed along four dimensions: text content completeness(TCC), image content completeness(ICC), image quality(IQ), and image-text synergy(ITS). Unlike traditional image–text consistency metrics, the image–text synergy (ITS) metric rewards tight, complementary alignment and penalizes redundancy between the textual and visual modalities.
To evaluate the effectiveness of the SEIR pipeline, we conduct human evaluation on both question and answer quality. Human evaluation results indicate that after three rounds of iteration, in terms of question generation, there is a 32% improvement compared to the initial version. In terms of answer generation, the improvements in the four dimensions compared to the initial answer are as follows: TCC (15%), ICC (11%), IQ (1%), and ITS (19%).
Similarly, we conducted experiments on the gap between several judgment models and human. The results show that the method based on the original MLLM has an average gap of 13% from humans, while SynJudge is closer to human preferences, with a deviation of only 5%.
Finally, we used InterSyn to conduct fine-tuning experiments on existing models. The experimental results show that the maximum improvements of the fine-tuned models in four capabilities are as follows: TCC (29.7%); ICC (10.3%); IQ (6%); ITS (52.1%).
Our main contributions are summarized as follows:
  • We present InterSyn, the first large-scale, multi-turn, instruction-following dataset for interleaved image-text question answering, covering a wide range of topics.
  • We propose SEIR, a method that ensures high-quality data generation across three refinement steps with minimal manual effort.
  • We introduce SynJudge, a multi-dimensional evaluation model for scoring interleaved outputs, enabling fine-grained assessment and effective feedback for model improvement.
  • We conduct comprehensive experiments demonstrating that InterSyn substantially enhances LMM performance in instruction alignment, image-text synergy, and multi-turn reasoning, contributing to the advancement of unified multimodal systems.
@@ 2 Related Work
Models for Interleaved Image-Text Generation. Recent advances in Multimodal Large Language Models (MLLMs), such as Flamingo [23], InternVL [24, 2], and Qwen-VL [3, 25], have substantially improved multimodal understanding. Meanwhile, diffusion models [26, 5, 27] achieve strong visual generation performance. To unify understanding and generation, models such as MiniGPT-5 [28] and Show-o [29] combine autoregressive text generation with diffusion-based image synthesis. More recent efforts  [6, 30, 31, 32] adopt unified autoregressive frameworks for interleaved generation. However, these models are not explicitly optimized for instruction-following and often struggle to maintain coherence and cross-modal consistency.
Figure 2: Overview of the InterSyn data construction pipeline.
Datasets for Interleaved Image-Text Generation.
High-quality interleaved image-text data is crucial for training multimodal models. Existing large-scale datasets like MMC4 [8], OBELICS [9], and CoMM [11] are primarily document-level corpora constructed from web sources, but often suffer from noise, weak alignment and low interaction intensity. Several benchmarks, such as OpenLEAF [18], InterleavedBench [22], and OpenING [19], focus on specific tasks. LeafInstruct [12] constructs an interleaved image-text dataset by filtering samples from existing corpora [8, 33, 14]. However, both benchmarks and datasets remain limited in scale and diversity.
Evaluation for Interleaved Image-Text Outputs Early multimodal evaluation metrics independently assessed text quality [34, 35] and image quality [36, 37]. Subsequent metrics [38, 39, 40, 41, 42] targeted image-text consistency, yet still inadequately evaluated the quality of interleaved outputs. More recent efforts, including InterleavedEval [12] and CoMM [11], leveraged large multimodal models (LMMs) for holistic assessment, but often exhibit misalignment with human judgment. OpenING [19] proposed IntJudge for pairwise comparisons, but it lacks fine-grained, quantitative scoring for individual responses, limiting its applicability for model training and refinement.
3 InterSyn Dataset
@@ 3.1 Overview
In this section, we describe the process of constructing the InterSyn dataset. Figure 2 illustrates the overall InterSyn construction pipeline, covering both the preparatory work and the SEIR method.
@@ 3.2 Dataset Preparatory Work
Question Filtering and Benchmark. We combined LLM-based filtering and expert review to select high-quality questions. Redundant, ambiguous, uncommon, and overly subjective samples were removed based on predefined criteria. In total, 500 diverse and high-quality questions—including both original and human-augmented samples—were selected to construct a benchmark.
Question Template Extraction. Based on the selected high-quality questions, we extracted generalized question templates to enable scalable question generation.
@@ 3.3 SEIR Method
The proposed SEIR method establishes a three-stage refinement process for generating high-quality multimodal data, as illustrated in Figure 2. Given a conversation turn parameter $T\in\mathbb{N}^{+}$ and a global refinement depth parameter $K\in\mathbb{N}^{+}$, for each turn $t\in\{1,\ldots,T\}$, the system generates data through a three-stage refinement process, which involves three cascaded modules: Question Refinement (QR), Answer Refinement (AR), and Image Refinement (IR). Each stage of refinement within a turn follows a Markovian property, where each iteration only depends on the immediate previous state. Across turns, topic consistency is maintained, and cross-turn coherence is achieved through contextual inheritance. This comprehensive approach allows for the generation of rich multimodal data with iterative optimization across multiple conversation turns and at each stage.
Let $\mathcal{T}$ denote the set of question templates, and $\mathcal{Z}$ denote the set of topics. For a conversation with $T$ turns, let the history up to turn $t-1$ be denoted as $\mathcal{H}^{(t-1)}$:
$\mathcal{H}^{(t-1)}=\{(q^{(i)},a^{(i)},i^{(i)})\}_{i=1}^{t-1}$ (1)
where each tuple $(q^{(i)},a^{(i)},i^{(i)})$ represents the question, text answer, and associated image at turn $i$. We define $\mathcal{H}^{(0)}=\emptyset$.
For each dataset sample, the data generation pipeline operates as follows:
[…]
The generation process for each conversation turn $t$ ($1\leq t\leq T$) proceeds as follows:
  1. 1.
    Turn Initialization: Sample a question template $\tau^{(t)}$ and generate an initial question $q_{0}^{(t)}$:
    $\tau^{(t)}\sim U(\mathcal{T})$ (3)
    $q_{0}^{(t)}=\mathcal{M}_{L}(p_{g}(\tau^{(t)},z,\mathcal{H}^{(t-1)}))$ (4)
    where $\mathcal{M}_{L}$ denotes the language model, and $p_{g}(\cdot)$ represents the generation prompt.
  2. 2.
    Question Refinement (QR): For each refinement iteration $k\in\{1,\ldots,K\}$:
    1. (a)
      Generate refinement suggestion:
      $s_{q}^{(t,k)}=\mathcal{M}_{L}(p_{s}(q_{k-1}^{(t)},z,\mathcal{H}^{(t-1)}))$ (5)
      where $p_{s}(\cdot)$ denotes the suggestion prompt, $q_{k-1}^{(t)}$ denotes the question at iteration $k-1$.
    2. (b)
      Refine question if $s_{q}^{(t,k)}\neq\emptyset$:
      $q_{k}^{(t)}=\mathcal{M}_{L}(p_{r}(q_{k-1}^{(t)},s_{q}^{(t,k)}))$ (6)
    where $p_{r}(\cdot)$ represents the refinement prompt. The iteration stops when $k=K$ or no further modifications are suggested ($s_{q}^{(t,k)}=\emptyset$), yielding the final question $q^{(t)}$.
  3. 3.
    Answer Refinement (AR): Generate initial answer and temp caption pair:
    $(a_{0}^{(t)},\gamma_{0}^{(t)})=\mathcal{M}_{L}(p_{g}(q^{(t)},\mathcal{H}^{(t-1)}))$ (7)
    where $a_{0}^{(t)}$ represents the initial answer and $\gamma_{0}^{(t)}$ represents the initial temp caption.
    For each refinement iteration $k\in\{1,\ldots,K\}$:
    1. (a)
      Generate refinement suggestion:
      $s_{a}^{(t,k)}=\mathcal{M}_{L}(p_{s}(a_{k-1}^{(t)},\gamma_{k-1}^{(t)},q^{(t)},\mathcal{H}^{(t-1)}))$ (8)
    2. (b)
      Refine answer and temp caption if $s_{a}^{(t,k)}\neq\emptyset$:
      $(a_{k}^{(t)},\gamma_{k}^{(t)})=\mathcal{M}_{L}(p_{r}(a_{k-1}^{(t)},\gamma_{k-1}^{(t)},s_{a}^{(t,k)}))$ (9)
    The process terminates when $k=K$ or no further refinements are suggested($s_{a}^{(t,k)}=\emptyset$), yielding the final answer $a^{(t)}$ and the temporary caption $\gamma^{(t)}$.
  4. 4.
    Image Refinement (IR): The temporary caption $\gamma^{(t)}$ is used as the initial caption $c_{0}^{(t)}$ to generate the initial image $I_{0}^{(t)}$:
    $c_{0}^{(t)}=\gamma^{(t)},\quad I_{0}^{(t)}=\mathcal{M}_{G}(c_{0}^{(t)})$ (10)
    where $\mathcal{M}_{G}$ is the text-to-image generation model.
    For each refinement iteration $k\in\{1,\ldots,K\}$:
    1. (a)
      Generate refinement suggestion for image caption:
      $s_{c}^{(t,k)}=\mathcal{M}_{V}(p_{s}(q^{(t)},a^{(t)},I_{k-1}^{(t)},\mathcal{H}^{(t-1)}))$ (11)
      where $\mathcal{M}_{V}$ denotes a VLM and $I_{k-1}^{(t)}$ represents the image at iteration $k-1$.
    2. (b)
      Refine caption if $s_{c}^{(t,k)}\neq\emptyset$:
      $c_{k}^{(t)}=\mathcal{M}_{L}(p_{r}(c_{k-1}^{(t)},s_{c}^{(t,k)}))$ (12)
    3. (c)
      Generate new image:
      $I_{k}^{(t)}=\mathcal{M}_{G}(c_{k}^{(t)})$ (13)
    The iteration stops when $k=K$ or no further refinements are suggested ($s_{c}^{(t,k)}=\emptyset$), yielding the final caption $c^{(t)}$ and the final image $I^{(t)}$.
3.4 Dataset Composition

### v2

# arXiv:2506.09427v2[cs.CV] 02 Mar 2026

@@ Abstract
Recent advancements in Large Multimodal Models (LMMs) have significantly improved multimodal understanding and generation. However, these models still struggle to generate tightly interleaved image-text outputs, primarily due to the limited scale, quality, and instructional richness of current training datasets. To address this, we introduce InterSyn, a dataset that features: (1) large scale, comprising 1.8M multimodal samples; (2) high quality, supported by our proposed Self-Evaluation with Iterative Refinement (SEIR) method for rigorous automated quality refinement; (3) rich instructional diversity, ensured through diverse well-designed question templates, based on human preferences and covering a 3500-topic hierarchy. These characteristics make InterSyn particularly well-suited for training LMMs in interactive image–text generation capabilities. To evaluate the capabilities, we propose SynJudge, a reliable automatic evaluator that aligns closely with human judge and outputs four interpretable scores: Text Content Completeness (TCC), Image Content Completeness (ICC), Image Quality (IQ), and Image–Text Synergy (ITS). These scores are complementary, covering both content and quality as well as cross-modal interaction, thereby forming a comprehensive evaluation framework. Experimental results on InterSyn subsets of up to 200K samples show that 25K–50K already yield substantial improvements, while scaling to 100K/200K brings further gains in TCC, ICC, and especially ITS, highlighting InterSyn’s: (1) scalability, as performance consistently improves with more data; (2) efficiency, as significant gains are achievable even with smaller subsets, making it accessible to researchers with varying computational resources.
@@ 1 Introduction
Multimodal understanding and generation are critical capabilities toward artificial general intelligence. In the past two years, multimodal large language models (MLLMs) (Liu et al., 2023; Chen et al., 2024c; Wang et al., 2024a) have shown remarkable performance in multimodal understanding and even surpassed humans in some areas, while we have also seen many impressive advances in high quality image generation (Esser et al., 2024b; Betker et al., 2023). However, these models are often limited to generating either text or image outputs in isolation, while real-world scenarios typically require tightly interleaved multimodal outputs.
@@
Recently, pioneer unified LMMs, such as Janus-Pro (Chen et al., 2025b), have shown great potential. However, they struggle to generate instruction-following interleaved image-text outputs, manifesting issues such as semantic drift, low image–text synergy, and poor image quality.
The main challenges lie in the limited scale, quality, and instructional richness of existing datasets. Even with existing datasets (Zhu et al., 2023; Laurençon et al., 2023; Chen et al., 2024a; b; Xu et al., 2024), these challenges remain due to their critical limitations: (1) Limited scale: Focus on narrow tasks and typically contain no more than tens of thousands of samples, limiting their applicability to broader real-world scenarios; (2) Unstable quality: Built on web-crawled sources (Yang et al., 2021; Laurençon et al., 2023) or reused corpora (Zhou et al., 2018; Zhu et al., 2023) with inconsistent quality and lack standardized quality control mechanisms; (3) Low interaction complexity: Rely on static documents or single-turn prompts, thus failing to capture the contextual richness and interleaved structure of authentic human conversations.
@@
To address the above issues, we introduce InterSyn—the first fully automated, high quality, large scale dataset for instruction-following, multi-turn question answering with interleaved image–text responses. With 1.8M single-turn and 50,000 multi-turn dialogues across 8 domains and 3,500 topics (as shown in Figure 1), InterSyn provides extensive coverage for diverse real-world scenarios. InterSyn’s quality is further enhanced by our Self-Evaluation with Iterative Refinement (SEIR) method, which embeds self-checking and feedback loops into each generation step, enhancing semantic completeness and cross-modal synergy. To ensure rich instructional diversity, we extract human-like query styles through question templates that capture varied linguistic structures, supported by a rich topic hierarchy for diverse, instruction-driven dialogues.
To evaluate interleaved image–text generation, several related benchmarks (An et al., 2024; Zhou et al., 2024; Chen et al., 2025a; Xia et al., 2024; Liu et al., 2024) have been proposed, but they still suffer from the following limitations: (1) Limited domain and scale: small, task-specific sets cannot cover realistic multi-turn dialogue needs; (2) Costly manual evaluation: accurate assessment still hinges on human, whose expense and delay hinder large scale, rapid benchmarking; (3) Weak alignment with human preference: current automatic metrics significantly diverge from human judgments on fine-grained multimodal reasoning; (4) Narrow evaluation scope: emphasise surface correctness while overlooking synergy and overall answer quality.
@@
Therefore, we propose SynJudge, a reliable and comprehensive judge model for evaluating interleaved image-text generation with high alignment to human judgment. SynJudge provides interpretable, quantitative feedback across four key dimensions: text content completeness (TCC), image content completeness (ICC), image quality (IQ), and image-text synergy (ITS). Unlike traditional image–text consistency metrics, the ITS metric focuses on rewarding tight, complementary alignment between the textual and visual modalities while penalizing redundancy.
To validate our contributions, we conduct a series of experiments. First, the effectiveness of our SEIR method is demonstrated, showing substantial quality improvements over a non-refined baseline. Additionally, comparisons with several existing models show that SEIR consistently outperforms them across all evaluation metrics. Next, we validate SynJudge, which exhibits the strongest alignment with human judgment, showing a significantly smaller deviation compared to zero-shot MLLM evaluators. Finally, we validate the utility of InterSyn by fine-tuning models on randomly sampled subsets, up to 200K examples. Results reveal significant performance gains with as few as 25K/50K samples, underscoring the dataset’s high data density and efficiency. These results confirm that our methods and dataset contribute to improved multimodal model performance.
Our main contributions are summarized as follows: (1) We present InterSyn, a large scale dataset of 1.8M high-fidelity samples, distinguished by its instructionally rich and complex dialogues that span over 3,500 topics; (2) We propose SEIR, a method that ensures high quality data generation across refinement steps with minimal manual effort; (3) We introduce SynJudge, a multi-dimensional evaluation model for scoring interleaved outputs, enabling fine-grained assessment and effective feedback for model improvement; (4) We conduct comprehensive experiments demonstrating that InterSyn substantially enhances LMM performance in instruction alignment, image-text synergy, and multi-turn reasoning, contributing to the advancement of unified multimodal systems.
@@ 2 Related Work
Models for Interleaved Image-Text Generation. Recent advances in MLLMs, such as Flamingo (Alayrac et al., 2022), InternVL (Chen et al., 2024d), and Qwen-VL (Wang et al., 2024a), have substantially improved multimodal understanding. Meanwhile, diffusion models (Ramesh et al., 2022; Betker et al., 2023; Esser et al., 2024a) achieve strong visual generation performance. To unify understanding and generation, models such as MiniGPT-5 (Zheng et al., 2023) and Show-o (Xie et al., 2024) combine autoregressive text generation with diffusion-based image synthesis. More recent efforts  (Team, 2024; Wu et al., 2024; Wang et al., 2024b; Chern et al., 2024) adopt unified autoregressive frameworks for interleaved generation. However, lacking targeted, high quality training data, these models are not explicitly optimized for instruction-following and often struggle to maintain coherence and cross-modal consistency—a gap InterSyn is specifically designed to fill.
Datasets for Interleaved Image-Text Generation. High quality interleaved image-text data is crucial for training multimodal models. Existing large scale datasets like MMC4 (Zhu et al., 2023), OBELICS (Laurençon et al., 2023), and CoMM (Chen et al., 2024b) are primarily document-level corpora constructed from web sources, but often suffer from noise, weak alignment and low interaction intensity. Several benchmarks, such as OpenLEAF (An et al., 2024), InterleavedBench (Liu et al., 2024), and OpenING (Zhou et al., 2024), focus on specific tasks. LeafInstruct (Xu et al., 2024) constructs an interleaved image-text dataset by filtering samples from existing corpora (Zhu et al., 2023; Huang et al., 2016; Zhou et al., 2018). However, both benchmarks and datasets remain limited in scale and instructional diversity. To this end, we introduce InterSyn, a large scale, high quality dataset with diverse, multi-turn dialogues and automated refinement.
Evaluation for Interleaved Image-Text Outputs Early multimodal evaluation metrics independently assessed text quality (Papineni et al., 2002; Lin, 2004) and image quality (Heusel et al., 2017; Salimans et al., 2016). Subsequent metrics (Hessel et al., 2021; Li et al., 2023; Lin et al., 2024; Chen et al., 2023; Lu et al., 2024) targeted image-text consistency, yet still inadequately evaluated the quality of interleaved outputs. More recent efforts, including InterleavedEval (Xu et al., 2024) and CoMM (Chen et al., 2024b), leveraged MLLMs for holistic assessment, but often exhibit misalignment with human judgment. OpenING (Zhou et al., 2024) proposed IntJudge for pairwise comparisons, but it lacks fine-grained, quantitative scoring for individual responses, limiting its applicability for model training and refinement. In contrast, the proposed SynJudge provides a more comprehensive evaluation by assessing both content completeness and modality synergy, aligning more closely with human judgment.
3 InterSyn Dataset and SynJudge Evaluator
@@ 3.1 Overview
In this section, we present a comprehensive framework for both the construction and evaluation of interleaved image-text generation. First, we describe the InterSyn dataset construction pipeline, detailing the preparatory work and our proposed SEIR method (illustrated in Figure 2). Subsequently, we introduce SynJudge, a specialized evaluator designed to assess the quality of interleaved outputs, along with the specific metrics and benchmarks established to validate the model’s performance.
@@ 3.2 Dataset Preparatory Work
Figure 2: Overview of the InterSyn Dataset Construction Framework. The top panel illustrates the dataset preparatory work, covering question collection, filtering, template standardization, and topic expansion. The bottom panel illustrates the Self-Evaluation with Iterative Refinement (SEIR) method, which employs a Generate-Evaluate-Refine loop across three cascaded stages. (1) Question Refinement (QR): An initial question $q_{0}$ is refined into the final question $q$ based on the topic $z$ and question template. (2) Answer Refinement (AR): Using $q$, the final answer $a$ and a temporary caption $\gamma$ are iteratively refined. (3) Image Refinement (IR): Initialize $c_{0}$ with $\gamma$. Refine the caption and image until the final image $I$ is produced. The right-side legend details the inputs and historical context ($\mathcal{H}^{(t-1)}$) used at each stage.
@@
Question Filtering and Benchmark. We combined LLM-based filtering and expert review to select high quality questions. Redundant, ambiguous, uncommon, and overly subjective samples were removed based on predefined criteria. In total, 500 questions were reserved as a basic benchmark.
Question Template Extraction. Based on the selected high quality questions, we constructed a set of generalized question templates that capture the conversational query style (i.e., the common linguistic forms people use to pose requests), independent of any specific knowledge content, thereby enabling scalable question generation. See the appendix §E for details.
@@ 3.3 SEIR Method
The Self-Evaluation with Iterative Refinement (SEIR) method enhances data quality via a Generate-Evaluate-Refine loop. For each conversation turn $t$, we define a generic refinement operator $\Phi$ that transforms an initial content (e.g., a question or answer) $x_{0}$ into a final content $x$ through $K$ iterations. The loop terminates when the evaluator returns no suggestions or when a maximum iteration depth $K$ is reached. The update rule at iteration $k$ is:
$x_{k}=\mathcal{M}_{refine}(x_{k-1},s_{k}),\quad\text{where }s_{k}=\mathcal{M}_{eval}(x_{k-1},\mathcal{C}).$ (1)
Here, $\mathcal{M}_{eval}$ acts as a judge, analyzing the current content $x_{k-1}$ against context $\mathcal{C}$ (e.g., the target topic, preceding question, or dialogue history) to provide specific suggestions $s_{k}$. $\mathcal{M}_{refine}$ then produces an improved version $x_{k}$. We apply this operator sequentially across three stages for the current turn $t$:
Step 1: Question Refinement (QR). We first generate an initial question $q_{0}^{(t)}$ from a template and topic. To ensure clarity and depth, we apply $\Phi$ to obtain the final question $q^{(t)}$:
$q^{(t)}=\Phi(q_{0}^{(t)}\mid\mathcal{C}=\{z,\mathcal{H}^{(t-1)}\}).$ (2)
The evaluator critiques the initial question $q_{0}^{(t)}$ based on the topic $z$ and conversation history $\mathcal{H}^{(t-1)}$, continuing the refinement until the question meets quality standards or the iteration limit is reached.
Step 2: Answer Refinement (AR). Using the final question $q^{(t)}$, we generate an initial answer-caption pair $(a_{0}^{(t)},\gamma_{0}^{(t)})$. The refinement loop ensures the text is comprehensive and the caption is relevant, yielding the final answer $a^{(t)}$ and a temporary caption $\gamma^{(t)}$:
$(a^{(t)},\gamma^{(t)})=\Phi((a_{0}^{(t)},\gamma_{0}^{(t)})\mid\mathcal{C}=\{q^{(t)},\mathcal{H}^{(t-1)}\}).$ (3)
Step 3: Image Refinement (IR). Finally, we generate and refine the image. We utilize the temporary caption $\gamma^{(t)}$ from AR as the initial caption $c_{0}^{(t)}$. In each iteration $k$, we generate an image $I_{k}^{(t)}$, which is then critiqued by a VLM ($\mathcal{V}$) against the textual context ($q^{(t)},a^{(t)}$) and dialogue history $\mathcal{H}^{(t-1)}$. The VLM’s feedback $s_{c}^{(k)}$ guides the generation of a refined caption $c_{k+1}^{(t)}$:
$I_{k}^{(t)}=\mathcal{G}(c_{k}^{(t)}),\quad s_{c}^{(k)}=\mathcal{V}(I_{k}^{(t)},q^{(t)},a^{(t)},\mathcal{H}^{(t-1)}),\quad c_{k+1}^{(t)}=\mathcal{M}_{refine}(c_{k}^{(t)},s_{c}^{(k)}).$ (4)
This cycle repeats until satisfaction or the iteration limit $K$ is met, yielding the final image $I^{(t)}$.
3.4 InterSyn Composition